OpenStudy (anonymous):

If f(x) = 3x+4, and g(x) = 1/(sqrt(x-2)), what is the explicit forms of f(g(x)) and g(f(x))?

OpenStudy (anonymous):

I think I have the answers, I just want to check with someone more experienced if they're correct. \[f(g(x)) = 3/\sqrt{x-2} + 4\] and \[g(f(x)) = 1/\sqrt{3x+2}\] ?

OpenStudy (eyust707):

Correct

OpenStudy (anonymous):

Thanks :)

OpenStudy (eyust707):

Great job giving a stab at it yourself!! thats the best way to learn